8 to the power of x=1/256
Answer by stanbon(75887)

8 to the power of x=1/256
-------------------------------
(2^3)^x = 1/(2^8)
2^(3x) = 2^(-8)
3x = -8
x = -8/3
===================
Cheers,
Stan H.